OpenStudy (anonymous):

3rd derivative of y=a/2(e^(x/a)+e^(-x/a)) (steps please)

OpenStudy (kainui):

Do you know what hyperbolic cosine and sine are?

OpenStudy (anonymous):

i have the 2nd derivative worked out already y'' = a/(2a^2 ) ( e〗^(x/a) + e^(-x/a) )

OpenStudy (anonymous):

needs to be worked out through differentiation

OpenStudy (kainui):

If you can get the second derivative, what's stopping you from getting the third?

OpenStudy (anonymous):

need to investigate (d^(n)y)/(dx^(n)) for all positive odd integers for n

OpenStudy (kainui):

Well, a useful identity to know is \[\cosh(t)=\frac{e^{t}+e^{-t}}{2}\]\[\sinh(t)=\frac{e^{t}-e^{-t}}{2}\] it's pretty straightforward to see that they are each their own derivatives, and what you have is essentially just the chain rule. For example, \[\frac{d^n}{dx^n}(Ae^{b\ x})=Ab^n e^{b\ x} \] should make some sense.

OpenStudy (anonymous):

thanks